Jefferson Lab HALL A Logbook for December 98 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

    
    

    Tuesday

  • 9361 12/29/98 11:46 jpchen water NMR: y=-2mm
  • 9360 12/29/98 11:40 jpchen Water test: HD2 problem
  • Monday

  • 9359 12/28/98 20:59 incerti Water cell moved down to -1 mm
  • 9358 12/28/98 19:16 bogdanw wire chamber gas system
  • 9357 12/28/98 18:49 bogdanw S1/S2 and Cherenkov
  • 9356 12/28/98 11:26 jpchen Water cell NMR study
  • 9355 12/28/98 11:16 jpchen HV's for detectors
  • Sunday

    
    

    Thursday

  • 9352 12/24/98 11:27 Seonho Choi Successful completion of E94-010 data taking
  • 9350 12/24/98 07:29 A. Deur restrected access
  • 9349 12/24/98 07:24 A. Deur ep measurement
  • 9347 12/24/98 06:31 A. Deur end of positron runs.
  • 9341 12/24/98 05:21 Roman Moller measurement
  • 9340 12/24/98 05:12 A. Deur moller done-->raster +helicity check
  • 9337 12/24/98 04:53 slifer target NMR
  • 9333 12/24/98 03:50 A. Deur run on empty target
  • 9330 12/24/98 03:26 A. Deur Back to beam permit
  • 9328 12/24/98 01:57 A. Deur Controled access
  • 9322 12/24/98 00:34 A. Deur Bscope + raster
  • Wednesday

  • 9319 12/23/98 23:44 K. McCormick ATTENTION: Problem with DAQ was in prescale.dat file
  • 9316 12/23/98 23:33 K. McCormick Problem with DAQ
  • 9315 12/23/98 23:31 K. McCormick Threshold Test
  • 9287 12/23/98 18:53 K. McCormick Series of NMR measurements made
  • 9286 12/23/98 18:46 K. McCormick Slit problem fixed
  • 9285 12/23/98 17:11 K. McCormick Picture to go with entry 9283
  • 9284 12/23/98 17:03 K. McCormick Problem with hall A slit in injector
  • 9283 12/23/98 17:02 K. McCormick Beam to Hall A during Target NMR?
  • 9280 12/23/98 16:15 K. McCormick HAPPEX charge assymetry
  • 9276 12/23/98 15:41 G Jones cuts at 2.059GeV
  • 9275 12/23/98 15:39 G Jones efficiency at 2.059Gev
  • 9260 12/23/98 12:07 G Jones cuts at 2.059GeV
  • 9259 12/23/98 12:05 G Jones efficiency at 2.059GeV
  • 9252 12/23/98 10:27 r gilman current calibration
  • 9249 12/23/98 10:12 S Incerti NMR
  • 9247 12/23/98 10:00 G Jones electron shower counter
  • 9243 12/23/98 09:23 G Jones phototube
  • 9242 12/23/98 09:08 G Jones efficiency at 2.059 GeV
  • 9241 12/23/98 09:06 G Jones cuts at 2.059 GeV
  • 9238 12/23/98 08:45 G Jones efficiency at 2.059GeV
  • 9237 12/23/98 08:43 G Jones cuts at 2.059GeV
  • 9234 12/23/98 07:30 A. Deur done with |_ @ 2324 MeV/c
  • 9233 12/23/98 07:26 P. Djawotho target polarization
  • 9226 12/23/98 04:25 A. Deur TS0 question
  • 9224 12/23/98 04:15 A. Deur // runs @ 2.423 GeV/c done
  • 9221 12/23/98 03:44 P. Djawotho target polarization
  • 9217 12/23/98 02:53 A. Deur TSO trouble
  • 9212 12/23/98 01:39 P. Djawotho target polarization
  • 9210 12/23/98 01:30 A. Deur spotsize
  • 9208 12/23/98 01:23 A. Deur Bscope
  • Tuesday

  • 9202 12/22/98 23:57 Seonho Choi trigger test
  • 9198 12/22/98 23:30 Wang Xu NMR
  • 9190 12/22/98 22:03 L.Todor Trigger tests
  • 9189 12/22/98 21:49 Wang Xu NMR &Laser 1
  • 9188 12/22/98 21:39 McCormick TS0 Rebooted
  • 9187 12/22/98 21:36 McCormick Current Calibration Postponed
  • 9174 12/22/98 19:17 Eugene Moller measurement
  • 9168 12/22/98 19:04 Eugene Moller measurement
  • 9164 12/22/98 18:44 Wang Xu NMR
  • 9163 12/22/98 16:03 F. sabatie No beam
  • 9154 12/22/98 14:16 incerti nmr
  • 9153 12/22/98 14:08 R. Michaels Trigger Latch TDC Fix
  • 9151 12/22/98 13:00 F. Sabatie Controlled Access
  • 9147 12/22/98 11:34 F. Sabatie Arc-en. during Run 3409
  • 9146 12/22/98 11:32 A. Saha Arc -energy 12/17/98
  • 9145 12/22/98 11:29 A. Saha Arc-energy measurement
  • 9144 12/22/98 11:12 F. Sabatie Run 3407
  • 9140 12/22/98 11:03 F. Sabatie Run 3407
  • 9134 12/22/98 08:20 A. Deur about ep.
  • 9131 12/22/98 07:52 A. Deur finished 2233 MeV/c kinematic
  • 9130 12/22/98 07:48 A. Deur adaqh3 rebooted
  • 9128 12/22/98 07:30 P. Djawotho target polarization
  • 9127 12/22/98 07:30 A. Deur beam back
  • 9125 12/22/98 06:46 A. Deur no beam
  • 9121 12/22/98 05:55 slifer show_cuts
  • 9118 12/22/98 04:45 P. Djawotho target polarization
  • 9116 12/22/98 04:25 A. Deur end of // runs @ 2.233 GeV/c
  • 9114 12/22/98 04:02 slifer show_cuts
  • 9106 12/22/98 00:52 P.Djawotho raster check
  • 9105 12/22/98 00:47 A. Deur Ask MCC for a Bscope
  • Monday

  • 9099 12/21/98 23:40 R. Michaels T4 deadtime effect
  • 9098 12/21/98 23:24 R. Michaels quick test suggestion
  • 9097 12/21/98 23:10 X. Jiang Trigger-2 and Trigger-4 of Run 3385
  • 9090 12/21/98 21:30 R.Michaels unknown deadtime corrections
  • 9085 12/21/98 19:08 Seonho Choi Cuts at 2.233GeV
  • 9084 12/21/98 19:07 Seonho Choi Efficiency at 2.233GeV
  • 9083 12/21/98 19:04 G Jones NMR measurement
  • 9082 12/21/98 18:58 Seonho Choi T2 and T4 rate reasonable
  • 9080 12/21/98 18:54 R.Michaels new trigger thresholds
  • 9079 12/21/98 18:47 Seonho Choi Threshold restored to 45mV
  • 9076 12/21/98 17:56 R.Michaels Problems of T2
  • 9071 12/21/98 15:55 O. Hansen Run 3385 is a test
  • 9064 12/21/98 14:02 R.Michaels T2 scaler discrepancy
  • 9062 12/21/98 13:05 O. Hansen NMR measurement
  • 9061 12/21/98 12:54 O. Hansen Finish p=2.846
  • 9059 12/21/98 11:56 O. Hansen DAQ trouble, ROC reboot
  • 9056 12/21/98 10:59 O. Hansen Beam returns
  • 9054 12/21/98 10:12 O. Hansen Beam gone
  • 9049 12/21/98 08:41 O. Hansen Argon gas restored
  • 9048 12/21/98 08:37 O. Hansen Status
  • 9044 12/21/98 07:51 slifer target NMR
  • 9043 12/21/98 07:49 A. Deur argon flow
  • 9040 12/21/98 06:53 slifer multiplicity of wires
  • 9038 12/21/98 06:50 A. Deur no argon flow
  • 9039 12/21/98 06:50 slifer VDC spectra
  • 9035 12/21/98 06:22 A. Deur gas alarm
  • 9032 12/21/98 04:22 slifer efficiency for 2.846 Gev/c
  • 9031 12/21/98 04:19 slifer cuts for 2.846
  • 9030 12/21/98 04:01 slifer target NMR
  • 9028 12/21/98 03:13 A. Deur end of perp run at 2625 MeV/c
  • 9026 12/21/98 02:09 A. Deur Bscope
  • 9023 12/21/98 00:40 Eugene Moller measurement
  • Sunday

  • 9019 12/20/98 23:58 Incerti nmr
  • 9007 12/20/98 17:04 Incerti beam back
  • 9004 12/20/98 16:00 G Jones NMR measurement
  • 9002 12/20/98 15:55 O. Hansen Pscaler results
  • 9001 12/20/98 15:32 O. Hansen Beam gone for "10 minutes"
  • 8999 12/20/98 15:04 O. Hansen planned repairs later today
  • 8996 12/20/98 14:53 O. Hansen Scintillator gains
  • 8995 12/20/98 14:52 O. Hansen More h-arm gas cer analysis
  • 8990 12/20/98 13:12 G. Jones NMR measurement
  • 8989 12/20/98 12:58 O. Hansen H-arm gas Cerenkov
  • 8988 12/20/98 12:49 O. Hansen Beam returns
  • 8986 12/20/98 12:48 O. Hansen Test run
  • 8983 12/20/98 11:20 O. Hansen Beam off
  • 8981 12/20/98 10:46 O. Hansen Hadron arm Cerenkov spectrum
  • 8978 12/20/98 10:27 R.Michaels trigger rates seem strange
  • 8974 12/20/98 09:38 O. Hansen Pedestal run
  • 8973 12/20/98 09:37 O. Hansen Ignore run 3356
  • 8968 12/20/98 08:28 slifer laser #6
  • 8966 12/20/98 08:22 slifer efficiency for 2.625 Gev
  • 8965 12/20/98 08:21 slifer cuts for new p
  • 8964 12/20/98 08:15 A. Deur Coda crash.
  • 8963 12/20/98 08:14 slifer pscaler asymmetry
  • 8958 12/20/98 07:52 A. Deur raster back.
  • 8957 12/20/98 07:49 A. Deur coda trouble for ep.
  • 8955 12/20/98 07:20 slifer NMR/ rotate back to //
  • 8954 12/20/98 07:15 A. Deur raster off.
  • 8953 12/20/98 07:13 A. Deur end of 3.086 GeV/c and ep
  • 8952 12/20/98 07:10 A. Deur ROC1 trouble
  • 8945 12/20/98 03:42 slifer target NMR
  • 8944 12/20/98 03:41 A. Deur Spotsize
  • 8941 12/20/98 02:12 slifer NMR /rotation to perpendicular
  • 8938 12/20/98 01:14 Wang Xu Deadtime_helicity_relationship
  • 8937 12/20/98 00:42 Wang Xu Helicity Check 2
  • 8936 12/20/98 00:35 Wang Xu Helicity Check 1
  • Saturday

  • 8905 12/19/98 17:50 Seonho Choi raster check
  • 8901 12/19/98 17:10 Eugene Moller measurement
  • 8900 12/19/98 17:09 E. Goldberg NMR
  • 8899 12/19/98 17:01 O. Hansen Bad beam conditions
  • 8898 12/19/98 16:54 O. Hansen 3He target in
  • 8896 12/19/98 15:51 G. Jones NMR measurement
  • 8895 12/19/98 15:46 O. Hansen Spot mystery solved
  • 8894 12/19/98 15:40 O. Hansen Spot check mysteries
  • 8889 12/19/98 15:13 O. Hansen Beam returns
  • 8888 12/19/98 14:35 Wang Xu Helicity Check 2
  • 8887 12/19/98 14:22 Wang Xu Helicity Check 1
  • 8885 12/19/98 13:22 G. Jones NMR Measurement
  • 8884 12/19/98 11:23 G. Jones NMR Measurement
  • 8881 12/19/98 10:35 O. Hansen HV display fixed
  • 8880 12/19/98 10:24 O. Hansen HV checks
  • 8879 12/19/98 10:03 O. Hansen Cosmics replay
  • 8878 12/19/98 09:59 O. Hansen Status update
  • 8877 12/19/98 09:24 G. Jones NMR measurement
  • 8874 12/19/98 07:17 Djawotho target polarization
  • 8873 12/19/98 06:55 R. Michaels E-arm and H-arm helicity flag check
  • 8872 12/19/98 06:37 R. Michaels switched to 1-TS DAQ
  • 8865 12/19/98 05:57 slifer MCC Update
  • Friday

  • 8864 12/18/98 22:02 kuss follow-up to 8861
  • 8863 12/18/98 21:43 kuss still no beam
  • 8862 12/18/98 21:40 kuss e94010_20621.dat.0
  • 8861 12/18/98 21:36 kuss e94010_20621.dat.0
  • 8858 12/18/98 16:57 E. Goldberg beam update
  • 8857 12/18/98 16:36 Mliang EPR, NMR calibraion
  • 8856 12/18/98 16:25 kuss still no beam
  • 8854 12/18/98 14:28 incerti mcc update
  • 8853 12/18/98 14:18 incerti nmr
  • 8852 12/18/98 13:41 incerti EPR Calibration History 2/2
  • 8851 12/18/98 13:38 incerti EPR Calibration History 1/2
  • 8850 12/18/98 13:23 incerti mcc update
  • 8849 12/18/98 11:20 incerti nmr
  • 8848 12/18/98 10:22 stevensm beamline vacuum problem
  • 8847 12/18/98 08:12 slifer shift summary
  • 8845 12/18/98 08:00 slifer beamline vaccum
  • 8844 12/18/98 07:49 P. Djawotho target polarization
  • 8843 12/18/98 07:27 slifer half-wave plate is in
  • 8842 12/18/98 07:23 h.fonvieille scin-kazu for runs 3324,20621
  • 8841 12/18/98 07:23 slifer pscaler asymmetry
  • 8840 12/18/98 07:02 P. Djawotho target polarization
  • 8839 12/18/98 06:56 slifer Beam down/beamline valves
  • 8838 12/18/98 06:40 slifer beam down
  • 8831 12/18/98 05:28 slifer beam down/NMR
  • 8830 12/18/98 05:25 slifer pscaler asymm
  • 8825 12/18/98 04:55 slifer happex
  • 8824 12/18/98 03:32 slifer 1st runs after down time
  • 8820 12/18/98 03:18 P. Djawotho target polarization
  • 8819 12/18/98 03:11 slifer cw beam
  • 8816 12/18/98 01:58 slifer beam down
  • 8815 12/18/98 01:56 slifer beam back then down
  • 8811 12/18/98 01:04 slifer pulsed beam to hall
  • 8809 12/18/98 00:00 Wang Xu Helicity Flag Check
  • Thursday

  • 8806 12/17/98 23:08 averett NMR measurement, beam estimate
  • 8803 12/17/98 21:16 averett NMR measurement
  • 8802 12/17/98 21:09 averett NMR measurement
  • 8801 12/17/98 20:54 averett Update from MCC
  • 8800 12/17/98 15:28 incerti NMR
  • 8798 12/17/98 14:20 Saha Energy Arc measurement
  • 8795 12/17/98 14:09 incerti once again temp drop
  • 8794 12/17/98 14:03 Saha Current Calibrations
  • 8790 12/17/98 13:46 incerti oven temp drop
  • 8782 12/17/98 12:08 incerti air flow adjustement
  • 8781 12/17/98 11:00 Choi-Papandr Charge asymmetry
  • 8765 12/17/98 09:50 Seonho Choi Heater tube problem fixed
  • 8764 12/17/98 08:43 Seonho Choi Target heater problems update
  • 8763 12/17/98 07:25 slifer target heater problems
  • 8762 12/17/98 01:42 Cates nmr epr calibration
  • Wednesday

  • 8760 12/16/98 16:20 Eugene Moller quad settings
  • 8758 12/16/98 11:43 Seonho Choi Maintenance Day starts Now
  • 8757 12/16/98 11:32 incerti NMR Spinup
  • 8756 12/16/98 11:12 incerti MAINTENANCE DAY
  • 8755 12/16/98 08:43 Seonho Choi Possible magnet problem, beam down.
  • 8753 12/16/98 05:52 slifer beam down
  • 8750 12/16/98 05:41 slifer happex
  • 8749 12/16/98 05:36 P. Djawotho target polarization
  • 8741 12/16/98 04:15 P. Djawotho efficiencies
  • 8740 12/16/98 04:10 h.fonvieille scin-kazu plots for run 3308
  • 8739 12/16/98 04:04 h.fonvieille scin-kazu plots for run 20609
  • 8734 12/16/98 03:13 P. Djawotho cuts on h-arm
  • 8733 12/16/98 02:38 slifer efficiency
  • 8732 12/16/98 02:28 P. Djawotho efficiencies for e-arm and h-arm
  • 8731 12/16/98 02:06 slifer beam trips
  • 8730 12/16/98 02:00 P. Djawotho cuts on e-arm
  • 8723 12/16/98 01:04 slifer raster
  • 8719 12/16/98 00:38 Eugene Moller measurement
  • 8718 12/16/98 00:33 h.fonvieille E arm dipole magnet power supply fault
  • Tuesday

  • 8713 12/15/98 23:18 Eugene Moller quad settings
  • 8712 12/15/98 22:56 Eugene Control access
  • 8711 12/15/98 22:02 R.Michaels HAPPEX asy (?)
  • 8710 12/15/98 21:35 averett charge asymmetry
  • 8705 12/15/98 21:10 averett NMR measurement
  • 8704 12/15/98 21:05 averett Beam arrives
  • 8699 12/15/98 19:02 averett NMR measurement, beam estimate
  • 8697 12/15/98 15:49 Pavlin Target polarization
  • 8696 12/15/98 12:23 Pavlin Target polarization
  • 8695 12/15/98 10:20 Pavlin Target polarization
  • 8694 12/15/98 08:34 Pavlin Target Polarization
  • 8693 12/15/98 06:22 P. Djawotho target polarization
  • 8692 12/15/98 04:09 P. Djawotho target polarization
  • 8691 12/15/98 02:09 P. Djawotho target polarization
  • 8690 12/15/98 01:58 P. Djawotho hadron arm limit
  • 8689 12/15/98 00:27 E. Goldberg target polarization
  • 8688 12/15/98 00:24 P. Djawotho momentum change
  • Monday

  • 8687 12/14/98 16:05 Pavlin Target polarization
  • 8686 12/14/98 14:32 Pavlin Air flow adjustment
  • 8685 12/14/98 14:16 Pavlin Target Polarization
  • 8684 12/14/98 14:15 Pavlin Target Polarization
  • 8683 12/14/98 13:04 R.Michaels HAPPEX DAQ diskquota problems
  • 8680 12/14/98 10:53 Pavlin Adjustment of air flow
  • 8679 12/14/98 10:32 Pavlin carbon Target removed
  • 8678 12/14/98 09:57 Pavlin Target polarization
  • 8667 12/14/98 07:11 A. Deur Happex run
  • 8659 12/14/98 06:13 A. Deur MCC current test
  • 8656 12/14/98 05:50 A. Deur kcoda
  • 8637 12/14/98 04:38 A. Deur Back to coll. IN
  • 8624 12/14/98 03:32 A. Deur start runs collimators out.
  • 8615 12/14/98 02:27 A. Deur Happex run
  • 8614 12/14/98 02:18 A. Deur changed pscales
  • Sunday

  • 8588 12/13/98 23:52 X. Jiang alignment
  • 8582 12/13/98 23:11 X. Jiang settled down for c12 run.
  • 8581 12/13/98 23:08 X. Jiang dead time for 4.9 micro Amp.
  • 8571 12/13/98 22:44 X. Jiang 2.3 micro Amp current.
  • 8568 12/13/98 22:36 Eugene Moller: new detector tests
  • 8565 12/13/98 22:33 X. Jiang rate too high. C12.
  • 8556 12/13/98 20:28 jpchen 12C target installed
  • 8555 12/13/98 19:20 X. Jiang run 3270
  • 8551 12/13/98 17:47 Eugene Moller measurement
  • 8550 12/13/98 16:33 jpchen half wave plate
  • 8540 12/13/98 15:10 T. Black beamspot check
  • 8537 12/13/98 14:55 T. Black NMR measurement
  • 8536 12/13/98 14:54 T. Black current calibration aborted
  • 8531 12/13/98 14:36 T. Black current calibration
  • 8522 12/13/98 13:04 van der Meer E-arm S1R1 is not dead
  • 8513 12/13/98 11:39 Mliang E-arm S1R1 PMT may be dead.
  • 8512 12/13/98 11:31 R. Michaels things to check with low thresh
  • 8511 12/13/98 11:26 van der Meer prescaler ps3=1
  • 8505 12/13/98 11:07 T. Black NMR measurement
  • 8502 12/13/98 10:33 R. Suleiman Hadron Field
  • 8500 12/13/98 09:56 T. Black NMR measurement
  • 8498 12/13/98 09:50 van der Meer changed scintilators thresholds to 45 mV
  • 8491 12/13/98 08:46 h.fonvieille question about rates T1/T2 and T3/T4
  • 8490 12/13/98 08:40 A. Deur ep not succesful
  • 8485 12/13/98 07:21 A. Deur Have to reboot adaqh3
  • 8478 12/13/98 06:59 A. Deur update
  • 8476 12/13/98 06:58 slifer Access/ NMR
  • 8474 12/13/98 05:12 slifer target NMR / Access
  • 8467 12/13/98 04:24 A. Deur total # of events for // runs
  • 8466 12/13/98 04:12 h.fonvieille E arm scintillators plots now
  • 8461 12/13/98 03:43 slifer target NMR
  • 8460 12/13/98 03:42 A. Deur begin anti // runs
  • 8459 12/13/98 03:41 A. Deur end of // runs.
  • 8457 12/13/98 03:17 h.fonvieille change of threshold in E arm scintillator
  • 8454 12/13/98 03:02 slifer current change
  • 8448 12/13/98 02:39 slifer scin ADC
  • 8447 12/13/98 02:36 A. Deur pedestal run.
  • 8445 12/13/98 02:19 A. Deur BSCOPE
  • 8439 12/13/98 02:09 slifer scintillator plane 1 electron
  • 8440 12/13/98 02:09 h.fonvieille more plots on E arm scintillator
  • 8438 12/13/98 01:58 h.fonvieille problem on E arm scintillator
  • 8431 12/13/98 01:17 slifer target NMR
  • 8426 12/13/98 01:00 slifer Scintillator KAZU
  • 8425 12/13/98 00:47 A. Deur Raster check
  • Saturday

  • 8401 12/12/98 23:07 Wang Xu NMR
  • 8399 12/12/98 23:02 K. McCormick BSCOPE for last two bpms
  • 8397 12/12/98 22:45 K. McCormick Changed EARM momentum to 2.273 GeV
  • 8381 12/12/98 21:05 K. McCormick High Voltage changed on HARM S2-4R
  • 8361 12/12/98 18:07 K. McCormick Runs 3219, 3220 Bad
  • 8360 12/12/98 18:05 K. McCormick Spectrometers moved to E: 17.492 and H:12.526 degrees
  • 8349 12/12/98 17:31 Wang Xu NMR
  • 8348 12/12/98 16:47 K. McCormick Changes to BCM screen Hac_bcm.AS.adl
  • 8347 12/12/98 16:16 BogdanW HA pmt replaced
  • 8344 12/12/98 14:54 T. Black access
  • 8340 12/12/98 14:09 T. Black Mac crash
  • 8337 12/12/98 14:03 T. Black 0.848 GeV data done
  • 8334 12/12/98 13:36 T. Black run 3214 is junk
  • 8331 12/12/98 13:03 van der Meer charge asymmetry
  • 8328 12/12/98 12:28 T. Black beamspot check
  • 8324 12/12/98 12:16 T. Black NMR measurement
  • 8321 12/12/98 11:56 Huber no end of run header (20544)
  • 8315 12/12/98 11:13 Rob van der changed scintillator threshold from 60mV to 30mV
  • 8298 12/12/98 09:46 Huber Scin_kazu plots
  • 8297 12/12/98 09:22 T. Black spot check
  • 8294 12/12/98 09:16 T. Black Momentu/polarization change
  • 8293 12/12/98 09:15 T. Black finished harp scan
  • 8292 12/12/98 09:13 Huber Harrp scan results
  • 8291 12/12/98 09:01 T. Black NMR measurement.
  • 8290 12/12/98 08:34 T. Black data acquisition at 1.059 GeV finished
  • 8278 12/12/98 07:05 P. Djawotho target polarization and field rotation
  • 8273 12/12/98 06:17 P. Djawotho magnet display problems
  • 8268 12/12/98 05:57 Adaq efficiency for run 20535
  • 8267 12/12/98 05:55 E. Cisbani Happex
  • 8266 12/12/98 05:51 Adaq efficiency for run 3201
  • 8265 12/12/98 05:39 P. Djawotho raster check
  • 8260 12/12/98 05:28 P. Djawotho beam is back!
  • 8257 12/12/98 04:12 P. Djawotho MCC update
  • 8256 12/12/98 02:56 P. Djawotho MCC sending pulsed beam
  • 8255 12/12/98 02:09 P. Djawotho MCC update
  • 8254 12/12/98 02:02 P. Djawotho target polarization
  • Friday

  • 8252 12/11/98 23:58 Brash Shift Summary
  • 8251 12/11/98 23:01 T. Black electronic Armaggedon
  • 8238 12/11/98 20:51 T. Black beamspot check
  • 8229 12/11/98 19:58 T. Black beam spot check
  • 8216 12/11/98 16:13 jpchen target, field rotation and NMR
  • 8214 12/11/98 16:10 gilman shift summary
  • 8215 12/11/98 16:10 Huber Harp Scan results
  • 8210 12/11/98 15:43 gilman charge asymmetries
  • 8200 12/11/98 14:43 Huber Show results plots for P0=0.977
  • 8199 12/11/98 14:41 gilman bscope
  • 8194 12/11/98 14:09 gilman spot size
  • 8193 12/11/98 14:06 gilman HAPPEX charge asymmetry
  • 8192 12/11/98 13:51 A. Deur Oven fixed, then NMR
  • 8187 12/11/98 13:30 gilman target problems
  • 8186 12/11/98 12:49 wangxu helicity flag check
  • 8185 12/11/98 12:00 R. Michaels Scintillator Trigger Efficiency
  • 8184 12/11/98 11:17 R. Michaels dhist multiplicity, etc
  • 8183 12/11/98 11:16 gilman VDC multiplicity problem
  • 8177 12/11/98 09:55 J. P. Chen Target polarization measurement
  • 8167 12/11/98 08:46 Huber Today's scin_kazu plots
  • 8153 12/11/98 07:14 h.fonvieille VDC multiplicity
  • 8150 12/11/98 07:04 h.fonvieille about vdc multiplicity
  • 8147 12/11/98 06:41 M.Iodice VDC's multiplicity spectra
  • 8144 12/11/98 06:32 h.fonvieille about glass windows seen by the spectrometers
  • 8141 12/11/98 06:12 P. Djawotho Q1 on electron arm
  • 8138 12/11/98 05:57 P. Djawotho oven heater problem?
  • 8135 12/11/98 05:37 h.fonvieille phi_target spectra E,H arms
  • 8131 12/11/98 05:01 h.fonvieille harp scan
  • 8129 12/11/98 04:53 h.fonvieille Arc Scan
  • 8120 12/11/98 02:50 P. Djawotho Field Rotation
  • 8119 12/11/98 02:39 h.fonvieille E-arm vcd spectra from offline gdh kumac
  • 8114 12/11/98 02:17 P. Djawotho Target Polarization
  • 8113 12/11/98 02:09 P. Djawotho B-scope software problems?
  • 8110 12/11/98 02:02 M.Iodice VDC E-arm wire map anomalies
  • 8109 12/11/98 01:57 P. Djawotho raster
  • 8106 12/11/98 01:38 P. Djawotho B-scope
  • 8105 12/11/98 01:37 P. Djawotho Happex Asymmetry
  • 8098 12/11/98 00:45 h.fonvieille junk run hadron arm
  • Thursday

  • 8089 12/10/98 23:53 Brash Shift Summary
  • 8084 12/10/98 23:05 Brash Charge Asymmetry
  • 8076 12/10/98 21:08 Adaq kumac updates
  • 8055 12/10/98 17:35 Brash Search Engine for HALOG
  • 8054 12/10/98 17:03 Adaq 20493 - End of Run failed, 20494 is bad.
  • 8046 12/10/98 16:06 Burtin Shift Summary
  • 8040 12/10/98 15:34 Saha Beam Energy for Dec 8
  • 8039 12/10/98 15:16 Burtin Now in Transverse configuration
  • 8038 12/10/98 15:15 Burtin Ptarget=36%
  • 8033 12/10/98 14:39 Burtin charge asymmetry
  • 8028 12/10/98 14:07 Huber show_results plots for P0=0.902
  • 8013 12/10/98 12:12 Rob van der End of run summaries missing
  • 8005 12/10/98 11:20 Huber power outage in CC
  • 7995 12/10/98 09:30 Huber Hadron Arm FP check
  • 7992 12/10/98 09:00 Huber Scin_kazu plots
  • 7987 12/10/98 08:19 burtin raster is OK
  • 7971 12/10/98 06:33 P. Djawotho target polarization
  • 7958 12/10/98 04:39 h.fonvieille show-results plot for electron arm
  • 7957 12/10/98 04:36 h.fonvieille show-results plot for hadron arm
  • 7956 12/10/98 04:28 P. Djawotho beam is back
  • 7953 12/10/98 04:20 P. Djawotho beam tuning
  • 7948 12/10/98 03:43 P. Djawotho Shift Summary
  • Wednesday

  • 7912 12/09/98 22:02 M.Iodice target polarization measurement
  • 7905 12/09/98 20:43 Eugene Sign of the beam polarization
  • 7903 12/09/98 20:34 Eugene Moller measurement
  • 7901 12/09/98 19:19 M.Iodice Moller measurement
  • 7898 12/09/98 19:15 Brash Modification to HALOG
  • 7889 12/09/98 17:56 Adaq - Bra Changes to Paw kumac files ....
  • 7883 12/09/98 17:09 M.Iodice half wave plate
  • 7872 12/09/98 15:32 black scintillator "fixed"
  • 7870 12/09/98 15:28 black beamspot for run 3122
  • 7869 12/09/98 15:27 black trying to fix a scintillator
  • 7862 12/09/98 14:30 Burtin Raster check
  • 7855 12/09/98 13:50 black junk run
  • 7853 12/09/98 13:48 Eugene Moller quad power supply
  • 7851 12/09/98 13:42 Eugene Moller measurement
  • 7850 12/09/98 12:40 E. Burtin Resume Moller measurement.
  • 7847 12/09/98 11:13 E. Burtin Helicity signal.
  • 7844 12/09/98 10:56 E. Burtin Controlled access
  • 7836 12/09/98 09:18 black beam current raised to 12 uA
  • 7835 12/09/98 08:49 black NMR measurement
  • 7832 12/09/98 08:32 prout beam spot
  • 7829 12/09/98 08:23 prout harp scan
  • 7825 12/09/98 08:10 prout Q3 hadron
  • 7822 12/09/98 07:02 prout MCC update
  • 7821 12/09/98 06:14 prout beam almost in hall
  • 7820 12/09/98 04:12 prout MCC update
  • 7819 12/09/98 02:39 prout water hose fixed
  • 7818 12/09/98 01:11 prout mcc update
  • 7817 12/09/98 01:02 prout prescalers
  • Tuesday

  • 7811 12/08/98 23:59 kuss no beam (still)
  • 7810 12/08/98 22:12 kuss CHL leak
  • 7809 12/08/98 21:35 kuss no beam
  • 7808 12/08/98 21:33 kuss HRSH Q3
  • 7807 12/08/98 21:20 kuss Q3 helium level failure
  • 7806 12/08/98 20:07 kuss no beam
  • 7805 12/08/98 20:00 A. Deur ep energy measurement done.
  • 7804 12/08/98 19:59 kuss no beam
  • 7803 12/08/98 17:42 kuss EP energy measurement
  • 7802 12/08/98 17:03 kuss beam energy measurement
  • 7800 12/08/98 15:38 R.Michaels finish of entry
  • 7799 12/08/98 15:06 R. Michaels downloaded the trigger, checked mlu
  • 7798 12/08/98 14:29 R. Michaels suggest to download trigger
  • 7792 12/08/98 11:49 T. Black runs 3107 and 20447 are junk
  • 7789 12/08/98 11:43 T. Black beam out of position
  • 7786 12/08/98 11:40 T. Black prescalers on triggers changed
  • 7782 12/08/98 11:22 wang xu Helicity flag
  • 7781 12/08/98 11:13 T. Black beam down, adjusting prescalers
  • 7780 12/08/98 10:40 Adaq beam straight through
  • 7777 12/08/98 10:19 Adaq Bscope done
  • 7754 12/08/98 06:20 A. Deur changed target configuration
  • 7747 12/08/98 05:01 A. Deur Q3 hadron tripped
  • 7735 12/08/98 03:35 A. Deur Raster
  • 7732 12/08/98 03:22 A. Deur Beam Back
  • 7731 12/08/98 03:21 Steffen NMR
  • Monday

  • 7728 12/07/98 17:45 Adaq bpm pedestal runs made
  • 7723 12/07/98 16:16 M. Iodice espace_kazu analysis
  • 7714 12/07/98 15:25 R. Michaels re: S1,S2 hit combinations
  • 7708 12/07/98 14:59 Adaq S1,S2 hit combinations
  • 7695 12/07/98 13:53 T. Black "non-intrusive" MCC test
  • 7684 12/07/98 12:32 J.Calarco coarse wire valleys
  • 7683 12/07/98 12:17 J.Calarco harpscan, bscope
  • 7682 12/07/98 12:14 T. Black polarization changed to parallel. NMR measurement made
  • 7681 12/07/98 12:05 R.Michaels re: coarse wire valleys
  • 7676 12/07/98 11:15 X. Jiang wire hits looks strange.
  • 7660 12/07/98 07:31 E. Cisbani raster size
  • 7651 12/07/98 05:50 A. Deur cavity trip
  • 7646 12/07/98 05:26 A. Deur change cut on e-arm
  • 7645 12/07/98 05:16 A. Deur Bscope
  • 7642 12/07/98 05:09 A. Deur Beam Back
  • 7641 12/07/98 05:05 A. Deur NMR
  • 7640 12/07/98 03:28 A. Deur beam down
  • 7637 12/07/98 03:23 A. Deur beam is back
  • 7634 12/07/98 02:05 A. Deur Beam still down.
  • 7633 12/07/98 01:50 A. Deur pedestal runs
  • 7632 12/07/98 01:32 A. Deur beam down
  • 7631 12/07/98 01:31 A. Deur NMR
  • 7630 12/07/98 01:29 A. Deur change kinematic (late entry)
  • 7629 12/07/98 01:28 A. Deur Bscope (late entry)
  • Sunday

  • 7617 12/06/98 23:35 Adaq - Bra Coarse VDC Spectrum Problem ?
  • 7604 12/06/98 20:30 X. Jiang spot check.
  • 7603 12/06/98 19:49 X. Jiang Target Polarization Direction Change.
  • 7602 12/06/98 19:48 X. Jiang Q3 Helium level of H-arm
  • 7597 12/06/98 19:07 R.Michaels spot check on ROC15 (E-arm)
  • 7592 12/06/98 18:15 X. Jiang Block 23 of h-arm lead-glass
  • 7591 12/06/98 18:07 X. Jiang H-arm vdc wir hits
  • 7590 12/06/98 17:59 X. Jiang Beam was off.
  • 7589 12/06/98 17:55 incerti NMR
  • 7588 12/06/98 17:38 Brash HALOG Problems
  • 7587 12/06/98 17:30 Adaq yet another test entry
  • 0 12/06/98 08:12 no_name no_key
  • 0 12/06/98 08:12 no_name no_key
  • 0 12/06/98 08:04 no_name no_key
  • 0 12/06/98 07:46 no_name no_key
  • 0 12/06/98 07:45 no_name no_key
  • 0 12/06/98 07:34 no_name no_key
  • 0 12/06/98 07:33 no_name no_key
  • 0 12/06/98 07:16 no_name no_key
  • 7530 12/06/98 05:35 Steffen NMR
  • 7529 12/06/98 05:34 A. Deur Beam is back
  • 7526 12/06/98 04:53 A. Deur beam back sooner than expected.
  • 7525 12/06/98 04:25 Steffen Spot Fixed
  • 7522 12/06/98 04:16 Steffen SPOT Problems
  • 7521 12/06/98 04:05 A. Deur Beam will stay down for awhile
  • 7520 12/06/98 04:03 A. Deur beam is down
  • 7519 12/06/98 03:48 A. Deur ep is working !
  • 7516 12/06/98 03:29 Steffen Spot Size
  • 7504 12/06/98 01:01 A. Deur Bscope
  • 7502 12/06/98 00:31 Steffen NMR
  • 7501 12/06/98 00:30 A. Deur Change kinematic.
  • Saturday

  • 7477 12/05/98 19:17 incerti nmr
  • 7467 12/05/98 17:07 incerti nmr
  • 7466 12/05/98 16:59 Adaq beam mask request
  • 7454 12/05/98 13:21 S.Malov Charge asymmetry
  • 7449 12/05/98 12:22 S.Malov Efficiency
  • 7445 12/05/98 12:11 S.Malov Beam spot
  • 7444 12/05/98 12:07 J. Calarco raster
  • 7443 12/05/98 12:02 J. Calarco harpscan
  • 7441 12/05/98 11:42 J. Calarco summary
  • 7415 12/05/98 06:47 P. Djawotho efficiency of electron arm
  • 7414 12/05/98 06:46 P. Djawotho dp for electron arm
  • 7413 12/05/98 06:44 P. Djawotho efficiency of hadron arm
  • 7412 12/05/98 06:43 P. Djawotho dp for hadron arm
  • 7411 12/05/98 06:35 P. Djawotho Target Polarization
  • 7402 12/05/98 05:00 P. Djawotho Target Polarization
  • 7393 12/05/98 03:02 Steffen HAPEX ASYM
  • 7388 12/05/98 01:38 Steffen RF Trip
  • 7381 12/05/98 00:31 P. Djawotho Target Polarization
  • 7380 12/05/98 00:22 slifer target NMR
  • 7379 12/05/98 00:17 slifer Large cuurent on target
  • Friday

  • 7366 12/04/98 23:38 S.Malov Harp scan
  • 7352 12/04/98 21:13 slifer Prescale 2,4
  • 7351 12/04/98 21:12 slifer Spot
  • 7346 12/04/98 21:02 slifer Electron shower
  • 7341 12/04/98 19:53 slifer NMR summ.
  • 7332 12/04/98 18:41 slifer beam halo testing
  • 7325 12/04/98 17:17 slifer beam position
  • 7324 12/04/98 17:17 slifer halo/system checks
  • 7323 12/04/98 17:09 slifer Target NMR
  • 7322 12/04/98 17:00 Adaq spot size
  • 7308 12/04/98 15:50 Eugene Target measurements
  • 7307 12/04/98 15:06 Eugene No Beam
  • 7306 12/04/98 14:55 Eugene Control access
  • 7305 12/04/98 14:21 Eugene Moller measurement
  • 7304 12/04/98 12:04 Eugene Mott measurement
  • 7303 12/04/98 12:03 Eugene End of control access
  • 7302 12/04/98 11:23 Eugene Control access
  • 7301 12/04/98 11:21 Eugene Start of Moller measurement
  • 7295 12/04/98 10:12 J. Gomez troubles setting electron dipole P0
  • 7289 12/04/98 09:03 Eugene End of 2.175 GeV/c parallel
  • 7287 12/04/98 08:59 Eugene Target going perpendicular
  • 7286 12/04/98 08:53 T.Black Target polarization measurement
  • 7279 12/04/98 07:12 Steffen Espace problem?
  • 7276 12/04/98 06:49 Steffen Beam Trips
  • 7271 12/04/98 05:30 Steffen NMR
  • 7258 12/04/98 02:52 mitchell DP spectrum at 2.175 GeV
  • 7257 12/04/98 02:14 Steffen NMR
  • 7256 12/04/98 02:13 Steffen Electron Dipole Problem
  • 7255 12/04/98 02:10 mitchell new e-arm cuts, h and e eff
  • 7250 12/04/98 01:15 Mitchell 2.357 kin completed
  • 7247 12/04/98 00:42 Ransome Disk filled up
  • Thursday

  • 7238 12/03/98 23:54 slifer B scope
  • 7237 12/03/98 23:53 slifer Shift summary
  • 7228 12/03/98 22:14 slifer random spot check
  • 7223 12/03/98 21:36 slifer target NMR
  • 7218 12/03/98 21:27 slifer Raster Amplitude change
  • 7204 12/03/98 18:56 slifer beam to hall
  • 7203 12/03/98 17:50 Saha Charge determination
  • 7202 12/03/98 17:34 Jack VDC card swap.
  • 7201 12/03/98 17:13 A. Deur ep fixed (?)
  • 7200 12/03/98 16:28 slifer target NMR
  • 7198 12/03/98 12:29 sabatie chl trip follow-up
  • 7197 12/03/98 12:09 gilman CHL trip may affect plans
  • 7194 12/03/98 11:55 gilman new BCM screen
  • 7190 12/03/98 11:37 R.Michaels ROC15 reboots (button works)
  • 7191 12/03/98 11:36 sabatie nmr
  • 7179 12/03/98 09:38 gilman new bpm screen
  • 7174 12/03/98 08:59 gilman big beam spot problem (?)
  • 7173 12/03/98 08:54 gilman beam spot off center of target
  • 7163 12/03/98 07:19 Steffen Switch to Perp.
  • 7159 12/03/98 07:09 Steffen NMR
  • 7155 12/03/98 06:45 Steffen Beam Position
  • 7150 12/03/98 05:31 Steffen Corrupt_Sync Fixed
  • 7148 12/03/98 05:28 Steffen ROC 1 Problem
  • 7147 12/03/98 05:22 Steffen Corrupt Sync
  • 7139 12/03/98 04:29 Steffen NMR
  • 7132 12/03/98 03:44 R.Michaels re: pscaler charge asy
  • 7131 12/03/98 03:32 R.Michaels reboots of ROC15
  • 7130 12/03/98 03:29 Steffen ROC 15 Restart
  • 7114 12/03/98 00:58 Steffen Raster Check
  • 7105 12/03/98 00:13 kuss bscope
  • Wednesday

  • 7103 12/02/98 23:48 slifer Target NMR
  • 7094 12/02/98 22:42 kuss charge asymmetry
  • 7085 12/02/98 21:17 slifer target NMR
  • 7070 12/02/98 19:31 kuss beam is back
  • 7068 12/02/98 19:04 kuss no beam
  • 7067 12/02/98 18:33 kuss no beam
  • 7066 12/02/98 18:24 slifer target NMR
  • 7065 12/02/98 17:50 kuss beam off
  • 7060 12/02/98 17:36 slifer spot
  • 7057 12/02/98 17:21 kuss pulsed beam
  • 7054 12/02/98 16:55 slifer beam trips/ halo
  • 7046 12/02/98 16:25 slifer spot
  • 7042 12/02/98 15:57 J. LeRose MCC: Beam Off
  • 7040 12/02/98 15:31 J. LeRose MCC: CREST
  • 7034 12/02/98 14:51 J. LeRose MCC: 10microA
  • 7031 12/02/98 14:47 J. LeRose MCC: Going to 10 microA
  • 7028 12/02/98 14:45 J. LeRose Moller Quad Wrong
  • 7025 12/02/98 14:24 incerti nmr
  • 7024 12/02/98 14:16 incerti epsilon factor
  • 7023 12/02/98 14:14 J.LeRose MCC: Beam off
  • 7018 12/02/98 13:55 J.LeRose MCC: Beam is Back
  • 7017 12/02/98 13:20 J.LeRose MCC: Beam Off
  • 7015 12/02/98 13:01 J.LeRose MCC
  • 7011 12/02/98 12:36 J.LeRose MCC: Beam Back
  • 7008 12/02/98 12:13 incerti NMR
  • 7007 12/02/98 12:11 J.LeRose MCC: beam interuption
  • 7006 12/02/98 12:07 Eugene Moller measurement
  • 7002 12/02/98 11:49 J.LeRose Harp Scan
  • 6998 12/02/98 10:53 X. Jiang H-arm, block #23.
  • 6997 12/02/98 10:51 J.LeRose BCM Calibration
  • 6995 12/02/98 10:39 J.LeRose BCM
  • 6994 12/02/98 10:24 J.LeRose Beginning Moller measurements
  • 6991 12/02/98 10:05 J.LeRose BCM & DAQ
  • 6988 12/02/98 09:21 J.LeRose BCM & DAQ problems
  • 6978 12/02/98 08:50 J.LeRose Run 20295 Problem
  • 6975 12/02/98 08:40 J.LeRose Downstream BCM Out
  • 6972 12/02/98 07:57 A. Deur NMR
  • 6963 12/02/98 06:25 A. Deur Spotsize
  • 6962 12/02/98 06:20 A. Deur NMR +switch to |_ config.
  • 6953 12/02/98 04:19 E. Burtin Beam current asymmetry
  • 6950 12/02/98 03:55 A. Deur New cuts on e arm.
  • 6949 12/02/98 03:48 A. Deur changed cut on h arm
  • 6939 12/02/98 03:05 A. Deur NMR
  • 6938 12/02/98 02:55 A. Deur spotsize
  • 6934 12/02/98 02:35 A. Deur beam back at two o'clock
  • 6933 12/02/98 02:33 A. Deur HARP scan
  • 6925 12/02/98 01:09 A. Deur Hall A access
  • Tuesday

  • 6919 12/01/98 23:22 slifer Effect of shields on NMR
  • 6918 12/01/98 21:49 slifer Effect of shields on NMR
  • 6917 12/01/98 18:56 slifer raster software problem
  • 6916 12/01/98 18:53 slifer target NMR
  • 6915 12/01/98 18:26 slifer beam permit
  • 6913 12/01/98 14:20 incerti nmr
  • 6912 12/01/98 12:02 incerti nmr
  • 6910 12/01/98 06:20 Steffen NMR Measurement
  • 6909 12/01/98 06:14 A. Deur ep target is not working.
  • 6908 12/01/98 05:23 Steffen Water Signal Fit
  • 6907 12/01/98 01:47 slifer target NMR